Dome closed for humidity at 8:05 pm. Reopened at 8:45 pm.
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2014-08-22 | 6.25 | 0.0 | 0.0 | 4.0 | 0.0 | 10.25 |
Links
[1] mailto:yara.jaffe@astro-udec.cl
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/1736/delete?destination=print/content/enr-ctio-blanco-4-m-2014-08-22
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/1736/edit?destination=print/content/enr-ctio-blanco-4-m-2014-08-22
[4] mailto:frieman@fnal.gov
[5] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/1737/delete?destination=print/content/enr-ctio-blanco-4-m-2014-08-22
[6] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/1737/edit?destination=print/content/enr-ctio-blanco-4-m-2014-08-22
[7] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/6183?destination=print/content/enr-ctio-blanco-4-m-2014-08-22
Real seeing was 4 arcsec in first quarter and 3 arcsec in second quarter. Not allowed to report on form.
DES Observing Team
Observers: Stephen Kent, Richard Kron
Telops: Hernan Tirado, Claudio Aguilera, Rodrigo Hernandez
Remote Support:
Observing Plan
Conditions
Accomplishments
Problems
Guider failed a few times because of very bad seeing
Notes for data management
Other